Rocks W & X are dated using an element with a half-life of 20 million years. Answer the following questions using the data obtained from these two rocks.

Rock W – Parent = 6.25% (1/16), Daughter = 93.75% (15/16)

Rock X – Parent = 25% (¼), Daughter = 75% (¾)

Calculate the age for each igneous rock. (Show your calculations.)

what would be a possible age for rock layer Z?
Answer #1

Following is the - complete Answer -&- Explanation: for the given: Question: in...typed & image format...

\RightarrowAnswer:

  1. Age of  Rock W = 80.0 million years
  2. Age of Rock X = 40.0 million years

\RightarrowExplanation:

Following is the complete: Explanation: for the above: Answer: in...image format...

Explanation : - Geven :- Half life 4 the elencat (marimetowa). tye - million your Purent Porentage tof mas Rock Wt memaining. formulas we can use the following formula in. l. 1 x (original mass of a after secay zn a cione tive element Where a numberUning tak for (Rock W), we will get the flowing: 2 s = 1 Age of Rock W ; ive. n = 4 ) = 4x + y = 4x (20 milion your ) = so miwe using : { Eg (1) for Rock x ² , we will get the folosing (25.0%) = 1 2 x ( 100-0%). Age t Rock X = nxty = 2* (20 milionye

Note: The details about Rock Z , hasn't been mentioned in the problem posted. Therefore, solving the part of the question: which deals with: Rock Z, can't be solved. - Thank You!